OpenStudy (anonymous):

(2n^4 3h^3)4? Medal to the best help :)

OpenStudy (kc_kennylau):

Laws of indices: \[\Large(ab)^n=a^nb^n\\\Large {\left(a^m\right)}^n=a^{mn}\]

OpenStudy (kc_kennylau):

Try to break the brackets first :)

OpenStudy (anonymous):

I messed up on the question (2n^4 3h^3) ^4

OpenStudy (skullpatrol):

The first thing you need is a question, what are you asked for?

OpenStudy (kc_kennylau):

For example, here is how you would simplify (3a^3 5b^4)^5: (3a^3 5b^4)^5 =(3a^3)^5 (5b^4)^5 =3^5 (a^3)^5 5^5 (b^4)^5 =243a^15 3125b^20

OpenStudy (kc_kennylau):

or in latex: \[ \Large(3a^3 5b^4)^5\\ \Large=(3a^3)^5 (5b^4)^5\\ \Large=3^5 (a^3)^5 5^5 (b^4)^5\\ \Large=243a^{15}3125b^{20} \]

OpenStudy (skullpatrol):

$$\Huge\text{Simplify}$$ $$\Huge(2n^4 3h^3)^4$$

OpenStudy (skullpatrol):

$$\Huge (ab)^m=a^mb^m$$

OpenStudy (skullpatrol):

$$\Huge (2n^4 3h^3)^4=2^4(n^4)^43^4(h^3)^4$$

OpenStudy (skullpatrol):

@j_jurgensturnt! are you still there?

OpenStudy (skullpatrol):

A mind is a terrible thing to waste